OpenStudy (anonymous):

Let y = (9 - 2 x)**6. Find the first and second derivatives of y. dy/dx = A ( 9 - 2x )B d2y/dx2= C ( 9 - 2x )D

OpenStudy (anonymous):

I mean y=(9-2x)^6

OpenStudy (y2o2):

\[y = (9 -2x)^6 \] \[{dy \over dx} = 6(9 -2x)^5(-2) = -12(9-2x)^5 \] \[{d^2y \over dx^2 } = -12(5)(9-2x)^4 (-2) = 120(9 - 2x)^4 \]